What are the wavelengths of electromagnetic wave in free space that have the following frequencies?

(a) 5.00 ✕ 1019 Hz
  pm
(b) 7.50 ✕ 109 Hz
cm
Answer #1

a) wavelength = speed of light / frequency

= c/f

= 3*10^8 / 5*10^19

= 0.6*10^-11 F

= 6 pm

(b) wavelength = 3*10^8 / 7.5*10^9

= 0.04 m = 4 cm